OpenStudy (anonymous):

Use common denominators to determine whether the ratios form a proportion. 6 : 9 and 9 : 12 A.Yes. They form a proportion. B.No. They do not form a proportion. please help

OpenStudy (anonymous):

Every proportion has a reduced form.

OpenStudy (anonymous):

\(6=2\times 3\). \(9=3\times 3\). \[ 2\times \cancel{3} : 3\times \cancel 3 \]So \[ 6:9 = 2:3 \]

OpenStudy (anonymous):

\(12 = 3\times 2\times 2\)\[ \cancel 3\times 3 :\cancel 3\times 2\times 2 \]So \[ 9:12 = 3:4 \]

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

B is correct.

OpenStudy (anonymous):

Yep so right nice thinking kid

OpenStudy (anonymous):

need help o never mind

OpenStudy (anonymous):

Use division to determine whether the ratios form a proportion. 1/4 and 6/24 A.Yes. They form a proportion. B.No. They do not form a proportion.

OpenStudy (anonymous):

Lol @wio can help u with theese questions I have know clue how to do them :p

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

Okay, so can you reduce 1/4? Can you reduce 6/24?

OpenStudy (anonymous):

nope and ya. to 1/6

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

wio?

OpenStudy (anonymous):

sorry, but \(6/24 \neq 1/6\).

OpenStudy (anonymous):

So \[ 6 = 2\times 3 \]And \[ 24 = 2\times 2\times 2\times 3 \]

OpenStudy (anonymous):

If you cancel, you get \[ \cancel{2\times 3}\to 1 \\ 2\times 2\times \cancel{2\times 3} \to 2\times 2 =4 \]